a sample of 12.8 liters of an ideal gas at 24.0 C and 745 torr is compressed and heated so that the volume is 7.2 liters and the temperature 72.0 C. What is the pressure in the container in torr?
Answer #1

From the ideal gas law PV\alpha T

\frac{P_1V_1}{T_1}= \frac{P_2V_2}{T_2}

Initial pressure, P_1 = 745 torr

Initial volume of an ideal gas, V_1 = 12.8 L

Initial temperature,T_1 = 24^ oC = 273+24 = 297 K

After compression and heating

Final volume, V_2 = 7.2 L

Final temperature,T_2 = 72^oC = 273+72 = 345 K

Final pressure in the container, P_2 = \frac{(P_1V_1)(T_2)}{(T_1)(V_2)}

\Rightarrow P_2= \frac{745 \times 12.8 \times 345}{297 \times 7.2} = 1538.5 torr